# Mirage Sources and Large TeV Halo-Pulsar Offsets: Exploring the Parameter Space

### v1

@@ III.3 Source separation and flux ratio
In this subsection, we discuss the separation distance (defined as the maximal separation distance between any two sources if two or more mirage sources are identified) and the flux ratio ($\chi$, defined as main source/mirage source, where the main source is the source with the smallest distance to the injection point and the mirage source is the farthest one) of the sources. For each chosen set of parameters, we carry out 100 runs and calculate the above two quantities of each run, and show their distribution in the plane of separation value and flux ratio. In all runs, the accelerator’s distance is fixed at $d=1\,$kpc. To reveal the difference more clearly, we set here a higher criterion TS$=25$ for the identification of a source. As such, the mirage ratio will decrease compared to the case shown in Figure 3 with the same parameters (i.e., not all runs can generate a mirage source at $d=1\,$kpc).
Our results are shown in Figure 4. In the top panel of the figure, we can see that the mean separation between the main source and the mirage source is positively correlated with $L_{\rm c}$. Given that the mirage source is formed due to the projection effect when particles propagate to another coherence of the magnetic field as discussed before, the separation value is correlated with the coherence length. On the other hand, we do not see a clear dependence of the flux ratio $\xi$ on $L_{\rm c}$.
@@
The regular-to-turbulent ratio $B_{\rm r}/B_{\rm t}$ also affects the source separation, although the impact is less important than the previous two parameters. We see from the bottom panel of Figure 4 that, when $B_{\rm r}/B_{\rm t}$ decreases, the separation distance (statistically) increases slightly. Adding a regular magnetic field is equivalent to adding a wave of infinitely long wavelength. As this regular component becomes stronger, the coherence becomes less distinct. Correspondingly, there will be less significant change of the mean field direction. In this sense, it is similar to the influence of the coherence length. Indeed, in the extreme case of $B_{\rm r}/B_{\rm t}\to\infty$, electrons just spiral around the regular magnetic field, so no mirage source and subsequently no source separation is expected.
@@ III.4 Offsets
Now we investigate the spatial offsets of the identified sources from the accelerator. We locate the source at a relatively large distance, 6 kpc, to reduce the mirage ratio, so that the definition of the offset is clear. Nevertheless, if more than one sources are identified, the offset is defined as the distance between the accelerator and the source closest to it.
For $L=L_{\rm ref}$, as shown in the left columns of Figure 5, there are only a few points sparsely distributed in the plane of the offset and the total counts of the source (as a representative of the source flux). This is owing to the large distance we chose, which results in a low flux of the source and hence not many of them are identified. For those with a source being identified, the obtained offsets are quite large for $L=L_{\rm ref}$. This is because when the source is faint and the distance is far, the position of the Gaussian centre is subject to the noise fluctuation and large PSF. If we assume a higher lumionsity, i.e., $L=10L_{\rm ref}$, as shown in the right column of Figure 5, the offsets become generally smaller. Meanwhile, as can be expected, the offset increases with larger $L_{c}$ and is inversely correlated to $B^{2}$. The reasons are similar to that for the separation distance, as discussed in the last section.
The $\gamma$-ray luminosity $\mathcal{L}$ of the source has an impact on source identification. This can be seen in the count maps displayed in Figure 6 and in Figure 7. The former shows a configuration of the turbulence with $L_{\rm c}=400$ pc, and the latter another configuration with $L_{\rm c}=4$ pc. In the configuration with $L_{\rm c}=400$ pc, the emission is found to be identified as two sources when the luminosity is equal to $10\mathcal{L}_{\rm ref}$ (see the left panel in Figure 6), while it can only be identified as one source when it is equal to $\mathcal{L}_{\rm ref}$ (see the right panel). In the configuration with $L_{\rm c}=4$ pc, the centre of the emission is not affected by the noise when the luminosity is set to $10\mathcal{L}_{\rm ref}$ (see the left panel in Figure 7), while it is strongly affected by the noise when it is set to $\mathcal{L}_{\rm ref}$ (see the right panel).
Figure 6: The counts map (relative intensity) for one configuration with $L_{\rm c}=400$ pc and distance of 6 kpc. The emission can be identified as two sources if the $\gamma$-ray luminosity is $\mathcal{L}=10\mathcal{L}_{\rm ref}$ (left panel), and it can only be identified as one source when $\mathcal{L}=\mathcal{L}_{\rm ref}$ (right panel). The noise has been removed.
Figure 7: The counts map (absolute counts number) for one configuration with $L_{\rm c}=4$ pc. The centre of the emission is not affected by the noise when $\mathcal{L}=10\mathcal{L}_{\rm ref}$ (left panel), and it is severely affected by the noise when $\mathcal{L}=\mathcal{L}_{\rm ref}$ (right panel). The noise is kept.
Table 1: Possible applicable sources
Source name PSR name Gl Gb Distance (kpc) $B_{\rm r}$ ($\mu$G) $B_{\rm t}$ ($\mu$G, $\beta=0.27$) Offset (pc) Offset () $\zeta$
1LHAASO J0216+4237u J0218+4232 139.508 -17.527 3.150 0.89 0.52 18.14 0.33 41%
1LHAASO J1837-0654u J1838-0655 25.246 -0.196 6.600 1.35 2.28 13.82 0.12 26%
1LHAASO J1912+1014u J1913+1011 44.485 -0.167 4.613 0.62 8.65 10.47 0.13 12%
1LHAASO J2005+3415 J2004+3429 71.425 1.571 10.783 0.25 0.85 47.05 0.25 84%
1LHAASO J2005+3050 J2006+3102 68.667 -0.530 6.035 0.27 1.85 21.07 0.20 10%
IV Possible applications
In the LHAASO catalog [18], some sources have been discovered in the vicinity of middle-aged pulsars, but the centres of these sources show significant offsets from the positions of the pulsars. See in particular those listed in Table 1. These offsets may be linked with the mirage halo phenomenon discussed here. We checked this possibility by simulating particle propagation around these pulsars, based on the Jansson-Farrar model of the Galactic magnetic field [30, 31]. This model provides the interstellar magnetic field strength, and the ratio between the regular and the turbulent magnetic field strengths at the position of each pulsar, which are important parameters in our simulations as discussed in the previous Section. We note that the turbulence level in this model needs be re-scaled by a factor $\beta$ to fit the boron-to-carbon ratio (B/C) [32, 22]. We find that $\beta=0.27$ gives a reasonable residence time of cosmic rays in our Galaxy by performing the same simulation as presented in Ref.[32], so that the predicted cosmic-ray grammage  [^1] The cosmic-ray grammage is given by $\langle X\rangle=1.4m_{\rm p}n_{\rm p}\langle t_{\rm res}\rangle c$, where $t_{\rm res}$ is the residence time, $n_{\rm p}$ is the hydrogen density in the ISM, $m_{p}$ is the mass of proton, and the prefactor 1.4 takes into account the contribution of nuclei (such as helium) in the ISM. is consistent with the cosmic-ray B/C measurement of AMS-02 [34]. In the simulation, we employ $L_{\rm c}=220\,$pc following the observation of the power spectrum of Galactic foreground rotation-measure maps [22]. For a smaller coherence length, $\beta$ could be smaller, but still at the level of $\sim 0.1$, see Ref. [32].
Then, we further assume an injection spectral index $\alpha=2$ for the electrons, and we simulate electron propagation around each of the pulsars that are mentioned above. To reduce the influence of the randomly generated turbulence on the results, we perform 100 separate runs of simulations for each pulsar and calculate the offset generated in each run. A threshold TS value of 49 (as is used in the LHAASO catalog), rather than 25, is used for source identification. The possibility of explaining the offsets with mirage halos is evaluated with $\zeta$, which is defined as the fraction of the runs showing a source offset larger than the observed one. The obtained values of $\zeta$ are shown in the last column of Table 1. We can see that relatively weak regular magnetic fields are present in the ISM around these pulsars, which is in favor of producing mirage halos. Indeed, the probability of showing large offsets are found to be above 10% for all the listed pulsars, demonstrating that mirage halos may be a reasonable explanation for these measured large offsets.
V Summary and Discussion
@@
Furthermore, based on the properties of identified sources, the influence of model parameters on the propagation of electrons have been evaluated with three indicators, namely, the mirage ratio, the maximum separation distance, and the offset. The mirage ratio is the proportion of runs with identification of two or more sources, and the maximum separation distance is the maximum distance between any two sources identified in one run. The offset is the distance between the pulsar’s position and the centroid of the closest identified source. Our main findings include:
  • The mirage ratio is negatively correlated with $d$, $B_{\rm r}/B_{\rm t}$ and $B^{2}$, and is positively correlated with $L_{\rm c}$ (strong dependence) and $\alpha$ (weak dependence). The mirage ratio is almost 100% at 1 kpc with the reference $\gamma$-ray luminosity. Even at a distance of 3 kpc, the mirage ratio could still reach more than 10% with reasonable combinations of parameters.
@@
  • The maximum separation is negatively correlated with $B^{2}$, and positively correlated with $L_{\rm c}$. The large maximum separation suggests that in the inter-arm region, and in the Galactic halo (where the coherence length is larger), two sources separated with 100 pc could still be linked with one another and belong to the same halo.
@@
  • The offset is also negatively correlated with $B^{2}$, and positively correlated with $L_{\rm c}$. Our calculations suggest that in the inter-arm region, and in the Galactic halo (where the coherence length is larger), the offsets between pulsar and the halo centroid could be as large as 50 pc.
@@
We have also discussed the possible application of this mechanism to explain some LHAASO sources with pulsar associations. For example, 1LHAASO J0216+4237u, J0212+4254u, and J0206+4302u could have a common origin, and be explained as mirage halos powered by the millisecond pulsar PSR J0218+4232. The large offsets between a few middle-aged pulsars and LHAASO sources may also be explained by our model naturally.
@@
In our simulations, we do not consider the feedback of relativistic particles on the magnetic field, such as the streaming instability. This simplification is reasonable for energetic electrons ($>100$ TeV), because the energy density of such high-energy electrons is usually very small compared with that of the interstellar medium, and hence the growth of such instabilities should be very slow.
@@
We also note that the synthetic turbulent magnetic fields used in our simulations are generated following Ref. [24], assuming an isotropic turbulence with Kolmogorov-type power spectrum. The nature of the turbulence in the interstellar medium may display some differences. In particular, studies have shown that Alfvénic turbulence is highly anisotropic at the gyro-scale of cosmic rays and thereby inefficient at scattering cosmic rays, while fast modes are isotropic and should play a dominant role in scattering cosmic rays [35], as supported by the recent analysis of radio polarisation observations [36]. Therefore, some differences must exist between the cosmic-ray scattering rates for particles propagating in our synthetic turbulence and that for particles propagating in magnetohydrodynamic turbulence. Although our synthetic turbulence may not be accurate on the cosmic-ray gyration scales, it still qualitatively captures the isotropic and filamentary nature of the interstellar turbulence on large scales, which is due to modes with large wavelengths. These modes are actually those which drive the filamentary propagation of cosmic-rays along magnetic field lines and thereby the formation of asymmetric gamma-ray emissions [21], mirage sources, and large offsets. Therefore, our calculations provide a reasonably good approximation of realistic interstellar turbulence, for the study of the appearance of asymmetric extended gamma-ray emissions around cosmic-ray sources. Furthermore, this method can take into account a wide range of length scales in the turbulence from the gyroradius of relevant electrons up to the turbulence injection scale with reasonable computing times. More realistic simulations, such as particle-in-cell simulations, or particle propagation in 3D cubes from magnetohydrodynamic simulations can, however, only cover a limited range of spatial scales, limiting the ranges of electron energies that can be studied with them.
@@ Acknowledgements.
We thank Chunkai Yu, Hao Zhou and Giovanni Morlino for discussion. This work is supported by NSFC under grants No. 12393852, 12350610239, 12393853, 12173018, 12121003, and U2031105.

### v2

@@ III.3 Source separation and flux ratio
In this subsection, we discuss the separation distance (defined as the maximal separation distance between any two sources if two or more mirage sources are identified) and the flux ratio ($\chi$, defined as main source/mirage source, where the main source is the source with the smallest distance to the injection point and the mirage source is the farthest one) of the sources. For each chosen set of parameters, we carry out 100 runs and calculate the above two quantities of each run, and show their distribution in the plane of separation value and flux ratio. In all runs, the accelerator’s distance is fixed at d=1d=1\,kpc. To reveal the difference more clearly, we set here a higher criterion TS=25=25 for the identification of a source. As such, the source number will decrease compared to the case shown in Figure 1 with the same parameters (i.e., not all runs can generate a mirage source at d=1d=1\,kpc).
Our results are shown in Figure 2. In the top panel of the figure, we can see that the mean separation between the main source and the mirage source is positively correlated with $L_{\rm c}$. Given that the mirage source is formed due to the projection effect when particles propagate to another coherence of the magnetic field as discussed before, the separation value is correlated with the coherence length. On the other hand, we do not see a clear dependence of the $N$ on $L_{\rm c}$.
@@
The regular-to-turbulent ratio $B_{\rm r}/B_{\rm t}$ exhibits a more complex influence on source separation, as shown in the bottom panel of Figure 2. In the case of pure turbulent field ($B_{\rm r}/B_{\rm t}=0$), electrons undergo random walk, generating numerous small-scale structures but lacking large-scale coherence, resulting in moderate separation distances. When $B_{\rm r}/B_{\rm t}\approx 1$, an optimal balance is achieved: the regular component provides directional guidance while maintaining sufficient turbulent complexity, leading to well-defined mirage sources with the largest separations. As $B_{\rm r}/B_{\rm t}$ increases to 3, the propagation becomes increasingly dominated by the regular field, suppressing the formation of secondary structures and reducing the separation. This non-monotonic behavior reflects the competition between field-line guidance and turbulent scattering in shaping electron distributions.
@@ III.4 Offsets
Now we investigate the spatial offsets of the identified sources from the accelerator. We locate the source at a relatively large distance, 1 kpc. Nevertheless, if more than one sources are identified, the offset is defined as the distance between the accelerator and the source closest to it.
The influence of $B_{\rm r}/B_{\rm t}$ on offsets follows a pattern similar to that for source separation. Pure turbulent fields ($B_{\rm r}/B_{\rm t}=0$) produce moderate offsets due to random scattering, while balanced configurations ($B_{\rm r}/B_{\rm t}\approx 1$) yield the largest offsets as electrons maintain directional coherence while exploring multiple magnetic coherence scales. Strong regular fields ($B_{\rm r}/B_{\rm t}=3$) confine electrons closer to the injection point, resulting in smaller offsets.
IV Summary and Discussion
@@
Furthermore, based on the properties of identified sources, the influence of model parameters on the propagation of electrons have been evaluated with three indicators, namely, the source number $N$, the maximum separation distance, and the offset. The maximum separation distance is the maximum distance between any two sources identified in one run. The offset is the distance between the pulsar’s position and the centroid of the closest identified source. Our main findings include:
  • The $N$ is negatively correlated with $d$, $L_{\rm c}$, and is positively correlated with $B^{2}$. The other dependences are relatively weak.
@@
  • The maximum separation is negatively correlated with $B^{2}$, and positively correlated with $L_{\rm c}$. It exhibits a peak at intermediate $B_{\rm r}/B_{\rm t}$ ratios, reflecting the competition between field guidance and turbulent scattering. The large maximum separation suggests that in the inter-arm region, and in the Galactic halo (where the coherence length is larger), two sources separated with 100 pc could still be linked with one another and belong to the same halo.
@@
  • The offset is negatively correlated with $B^{2}$, and positively correlated with $L_{\rm c}$. Similar to separation, offsets are maximized at balanced $B_{\rm r}/B_{\rm t}$ ratios. Our calculations suggest that in the inter-arm region, and in the Galactic halo (where the coherence length is larger), the offsets between pulsar and the halo centroid could be as large as 50 pc.
@@
We have also discussed the possible application of this mechanism to explain some LHAASO sources with pulsar associations. For example, 1LHAASO J0216+4237u, J0212+4254u, and J0206+4302u [30] could have a common origin, and be explained as mirage halos powered by the millisecond pulsar PSR J0218+4232. The large offsets between a few middle-aged pulsars and LHAASO sources may also be explained by our model naturally.
@@
In our simulations, we do not consider the feedback of relativistic particles on the magnetic field, such as the streaming instability. This simplification is reasonable for energetic electrons (>100>100 TeV), because the energy density of such high-energy electrons is usually very small compared with that of the interstellar medium, and hence the growth of such instabilities should be very slow.
@@
We also note that the synthetic turbulent magnetic fields used in our simulations are generated following Ref. [24], assuming an isotropic turbulence with Kolmogorov-type power spectrum. The nature of the turbulence in the interstellar medium may display some differences. In particular, studies have shown that Alfvénic turbulence is highly anisotropic at the gyro-scale of cosmic rays and thereby inefficient at scattering cosmic rays, while fast modes are isotropic and should play a dominant role in scattering cosmic rays [31], as supported by the recent analysis of radio polarisation observations [32]. Therefore, some differences must exist between the cosmic-ray scattering rates for particles propagating in our synthetic turbulence and that for particles propagating in magnetohydrodynamic turbulence. Although our synthetic turbulence may not be accurate on the cosmic-ray gyration scales, it still qualitatively captures the isotropic and filamentary nature of the interstellar turbulence on large scales, which is due to modes with large wavelengths. These modes are actually those which drive the filamentary propagation of cosmic-rays along magnetic field lines and thereby the formation of asymmetric gamma-ray emissions [21, 33], mirage sources, and large offsets. Therefore, our calculations provide a reasonably good approximation of realistic interstellar turbulence, for the study of the appearance of asymmetric extended gamma-ray emissions around cosmic-ray sources. Furthermore, this method can take into account a wide range of length scales in the turbulence from the gyroradius of relevant electrons up to the turbulence injection scale with reasonable computing times. More realistic simulations, such as particle-in-cell simulations, or particle propagation in 3D cubes from magnetohydrodynamic simulations can, however, only cover a limited range of spatial scales, limiting the ranges of electron energies that can be studied with them.
@@ Acknowledgements.
We thank Chunkai Yu, Hao Zhou and Giovanni Morlino for discussion. This work is supported by NSFC under grants No. 12393852, 12350610239, 12393853, 12173018, 12121003, and U2031105, by the Fundamental Research Funds for the Central Universities under No. 020114380057, and by K. C. Wong Educational Foundation.